OpenStudy (anonymous):

Simplify this equation if possible: 7 y+4=5y-2

OpenStudy (anonymous):

So this is where I got at: 7y+4=5y-2 -5y -5y --------------- So that equals 2. But, I get confused on what to do next.

jimthompson5910 (jim_thompson5910):

you should have 2y + 4 = -2

jimthompson5910 (jim_thompson5910):

subtract 4 from both sides to get 2y + 4 = -2 2y + 4-4 = -2-4 2y + 0 = -6 2y = -6

OpenStudy (anonymous):

Oh. okay that makes sense! So: 2y + 4 = -2 -4 -4 --------------- -6/2y= -3

jimthompson5910 (jim_thompson5910):

good, y = -3 is your final answer

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

yw